# Market-Driven Equilibria for Distributed Solar Panel Investment

### v1

@@ Appendix A Proofs for Section IV
Appendix B Proofs for section V
B-A Proof of Lemma 4 and Proposition 1
@@ B-B Proof of Theorem 1
We defer the detailed proofs of the first three steps to Appendix C, and summarize their results below in the form of lemmas.
@@
Appendix C Proof of Lemma 5 and Lemma 6
C-A Proof of Lemma 5
@@
which can be applied in (C-A) to reformulate it as
@@
Adding inequalities (53) and (55), and letting $m=m_{1}+m_{2}$ and $M=M_{1}+M_{2}$, we obtain (C-A), which in turn implies (50).
@@ C-B Proof of Lemma 6
Using this result in (C-B), we can reformulate (C-B) as
@@
Appendix D Proofs for Section VI
D-A Proof of Proposition 2
@@
$\max_{c\in\mathbb{R}_{+}}\enspace\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\Biggr{[}$ $L_{t}\mathbb{1}\!\left\{cG_{t}\leq L_{t}\right\}\!\int_{0}^{F_{V}^{-1}\left(1-\frac{cG_{t}}{L_{t}}\right)}v_{i}\,\,\mathrm{d}F_{V}(v_{i})$
@@ D-B Proof of Theorem 2
The steps follow the same structure as in the proof of Theorem 1. Therefore, we begin by introducing new lemmas analogous to Lemmas 5 and 6.
@@
Appendix E Proof of Lemma 8 and Lemma 9
E-A Proof of Lemma 8

### v2

@@ Appendix B Proofs for Section IV
Appendix C Proofs for section V
C-A Proof of Lemma 4 and Proposition 1
@@ C-B Proof of Theorem 1
We defer the detailed proofs of the first three steps to Appendix D, and summarize their results below in the form of lemmas.
@@
Appendix D Proof of Lemma 7 and Lemma 8
D-A Proof of Lemma 7
@@
which can be applied in (D-A) to reformulate it as
@@
Adding inequalities (65) and (67), and letting $m=m_{1}+m_{2}$ and $M=M_{1}+M_{2}$, we obtain (D-A), which in turn implies (62).
@@ D-B Proof of Lemma 8
Using this result in (D-B), we can reformulate (D-B) as
@@
Appendix E Proofs for Section VI
E-A Proof of Proposition 2
@@
$\max_{c\in\mathbb{R}_{+}}\kern 5.0pt\sum_{t\in\mathcal{T}}\frac{\widetilde{T}}{T}\mathbb{E}\Biggr[$ $L_{t}\mathbb{1}\!\left\{cG_{t}\leq L_{t}\right\}\!\int_{0}^{F_{V}^{-1}\left(1-\frac{cG_{t}}{L_{t}}\right)}v_{i}\,\,\mathrm{d}F_{V}(v_{i})$
@@ E-B Proof of Theorem 2
The steps follow the same structure as in the proof of Theorem 1. Therefore, we begin by introducing new lemmas analogous to Lemmas 7 and 8.
@@
Appendix F Proof of Lemma 10 and Lemma 11
F-A Proof of Lemma 10